Solve the following system of linear equation by the substitution method.
y=3x-4
x+y=32

Answer by checkley71(8403)

y=3x-4 then x+(3x-4)=32 or x+3x-4=32 or 4x=32+4 or 4x=36 or x=9 then y=3*9-4 or y=27-4 or y=23